# version 1.0

### v6

@@ 2.2.3 Quantitative Definitions of Stabilizability
The quantified notion of stabilizability we introduce below is more refined than driving the system to zero. We require the ability to drive the system to zero under adversarial bounded noise in the dynamics. In classical control theory this is called BIBO-stability, which stands for bounded-input-bounded-output.
This definition can apply to any dynamical system and is an important concept in nonstochastic control.
@@
A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if and only if the following holds. For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$ for all $t\in[T]$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}),\mathbf{w}_{t}).$
Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, that is,
$\forall t\ ,\ \|\mathbf{x}_{t}^{\pi}\|\leq\gamma.$
@@
2.3 Intractability of Equilibrium, Stabilizability, and Controllability
Before formally proving the computational hardness of determining various solution concepts of dynamical systems, we start with a few intuitive and convincing examples of the complexities that arise. Consider the following dynamical system over the complex numbers:
@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
The Julia set is now defined as
$J_{c}=\{z\in C\ \colon\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
where $f^{t}_{c}(x)=f_{c}(f_{c}(\dots f_{c}(x)))$ denotes $t$ successive applications of the function $f_{c}$. In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (that is not $\infty$). The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, that is, there is no Turing machine that can decide it in finite time! (See e.g. [BY08].) The complexity that arises near the boundary of the Julia set is depicted in Figure 2.7.
Another compelling visual illustration of the complexity of equilibrium in dynamical systems is the Lorenz attractor. This is a strange equilibrium resulting from simple dynamics that can be described using the following three equations over three variables, called the Lorenz equations [Lor63],
$x_{t+1}$ $=\sigma(y_{t}-x_{t})$
$y_{t+1}$ $=x_{t}(\rho-z_{t})-y_{t}$
$z_{t+1}$ $=x_{t}y_{t}-\beta z_{t}.$
The three dimensional (3D) trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly, as depicted in Figure 2.7. For formal properties of this attractor set, see [Str14, Tuc99]. This shows that even studying a solution concept itself is a formidable task.
For a formal proof of attractiveness in the continuous case, see section 9.2 of [Str14] or the original paper of Lorenz [Lor63].
Figure 2.6: Julia set.
@@ 2.3.1 Computational Hardness of Stabilizability
Consider the following problem: Given a dynamical system, can we determine if it stabilizable, that is, for every possible starting state, does there exist a choice of control inputs that drives the system to the null state? We now show that this decision problem is computationally hard.
@@
For a family of dynamical systems described as polynomials with integer coefficients, determining the stabilizability of any member is NP-hard.
We show that this problem can be reduced to the 3SAT problem, a well-known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
@@
We first claim that checking if a given polynomial has a zero is NP-hard. This is achieved by a reduction from 3SAT. Recall the 3SAT problem: Given a Boolean formula in the conjunctive normal form over $n$ variables and with $m$ clauses where each clause is restricted to three literals, such as
$F(x_{1:n})=(x_{1}\lor\neg x_{2}\lor x_{3})\land(\neg x_{1}\lor\neg x_{2}\lor x_{3})\land(x_{1}\lor x_{3}\lor x_{4}),$
decide if there exists a satisfying Boolean assignment.
Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-three polynomial as
$g_{x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow(1-u_{1})\cdot(1-u_{2})\cdot(1-u_{3})$
$g_{\neg x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot(1-u_{2})\cdot(1-u_{3})$
$g_{\neg x_{i}\lor\neg x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot(1-u_{3})$
$g_{\neg x_{i}\lor\neg x_{j}\lor\neg x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot u_{3},$
and the degree-six polynomial for the entire formula as
$p_{F}(\mathbf{u})=\sum_{i=1}^{n}(u_{i}(1-u_{i}))^{2}+\sum_{j=1}^{m}(g_{c_{j}}(\mathbf{u}))^{2}.$
For a disjunctive clause to be false, all its literals must evaluate to false. Therefore, given any satisfying Boolean assignment $x_{1:n}$ for the formula $F$, it holds that $p_{F}(x_{1:n})=0$.
Similarly, given any $\mathbf{u}\in{\mathbb{R}}^{n}$ such that $p_{F}(\mathbf{u})=0$, it must be that $\mathbf{u}_{i}\in\{0,1\}$ for all $i$. Furthermore, since $p_{F}$ is a sum of squares, the Boolean assignment $x_{1:n}=\mathbf{u}$ must satisfy every clause in $F$. This shows the hardness of determining the existence of a zero for polynomials.
Now, consider the scalar-state dynamical equation given by
$x_{t+1}=(1-\varepsilon)x_{t}+p_{F}(\mathbf{u}_{t}),$
with $p$ being the polynomial family we just constructed and $\varepsilon>0$ an arbitrarily small constant. For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which as we just proved is NP-hard. ∎
The proof establishes that even for a dynamical system that is a polynomial of degree six, checking for stabilizability is NP-hard. Exercise 2.5 asks the reader to strengthen this result and show that checking the stabilizability of dynamical systems describable as polynomials of degree four is also NP-hard.
@@ 2.4 Bibliographic Remarks
The Lorenz equations were introduced to model the phenomenon of atmospheric convection [Lor63]. For detailed information on the mathematical properties of the Lorenz attractor, see [Str14, Tuc99]. For a rigorous study of Julia sets and the Mandelbrot set, see [BY08].
Nonlinear dynamical systems admit a variety of stability criteria; we chose to define a few that are useful for further developments in the text. A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thus obviating the need for a trajectory-based analysis. An interested reader might refer to [Kha15] for an expanded discussion on how these stability criteria compare.
An excellent survey on the computational complexity of control is [BT00]. For a more recent and stronger result on stabilizability, see [Ahm12]. The NP-hardness theorem in this chapter appeared in [Ahm16]. The computational complexity of mathematical programming – both constrained and unconstrained – is commented upon in [Haz19].
@@ 2.5 Exercises
2.1. Strengthen Theorem 2.5 to show that checking whether a dynamical system that is a polynomial of degree four is stabilizable is NP-hard. Hint: Reduce from the NP-hard problem of MAX-2-SAT instead of 3-SAT.
@@
$\mathbf{x}_{t+1}=\frac{A\mathbf{x}_{t}}{\|\mathbf{x}_{t}\|}.$
What are the stable and attractive equilibrium points $\mathbf{x}\in\mathcal{R}^{d}$ of the system? Are they also globally attractive? Your answer may depend on $A$.
2.3. Let $\alpha,\beta,\gamma>0$. Does the SIR-model, as is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)? What about controllability? We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ must hold at all times.
@@ 2.4. Consider the double integrator system:
The previous chapters described and motivated the problem of control in dynamical systems. We have seen that in full generality, even simple questions about dynamical systems such as stabilizability and controllability are computationally hard. In this chapter, we consider a different representation of the control problem for which computationally efficient methods exist.
This chapter introduces a setting of equal generality as that of the control problem from the previous chapter: reinforcement learning in Markov Decision Processes (MDP). However, the representation of states and dynamics in an MDP is more explicit and verbose. In such a case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming and the accompanying characterization of optimal policies via the Bellman equation.
@@ 3.1 Reinforcement Learning
Reinforcement learning is a subfield of machine learning that formally models the setting of learning through interaction in a reactive environment. This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling. In reinforcement learning, we have an agent and an environment. The agent observes its position (or state) in the environment and takes actions that transition it to a new state. The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
Figure 3.1: The agent-environment interaction loop in reinforcement learning.
Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment. This methodology has led to some notable successes: machines have learned how to play Atari games and how to beat the highest-rated human players at Go.
A crucial assumption for the reinforcement learning model is as follows:
The Markovian assumption: Future states, observations, and rewards are independent of the past given the current state.
The reader can imagine situations where such an assumption is reasonable. For example, in the inverted pendulum setting described in Section 2.1.3, the assumption is clearly valid. It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
Other problems may call into question this assumption. For example, in the medical ventilation problem, it may very well be important how a certain state was reached, as future treatment may depend on the medical history of the patient.
It can be argued that the Markovian assumption can always be made to work by adding more features to the state. In extreme cases, the state can contain the entire history, which makes the assumption true by default. However, this makes the state representation inefficient.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom left square, and its goal is to reach the +1+1 reward. The blocks +1,1+1,-1 represent terminal states, which means that the game ends when the robot reaches them. The block ‘X’ represents an obstacle through which the robot cannot travel.
@@
In this example, the robot has controls that allow it to move left / right / forward / backward. However, the controls are not reliable and, with $0.8$ probability, the robot moves wherever it intends to. With the remaining probability $0.2$, the robot moves sideways, that is, with probability $0.1$ to each side.
Notice that the Markovian assumption clearly holds for this simple example. What is an optimal solution to this example? A solution concept for a reinforcement learning problem is a mapping from state to action. This mapping is called a “policy”, and depends both on the structure of the problem and on the rewards.
For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent +1+1 reward, as given in figure 3.3.
@@
Figure 3.3: The optimal policy (mapping from state to action) with reward for each step taken being -0.04.
Conversely, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in Figure 3.4.
@@
Figure 3.4: Here, we change the reward of each step to be -2. The optimal policy changes: now the robot is more concerned with ending the game as quickly as possible rather than trying to get to the +1 reward.
We next describe a formal model to capture this and any other reinforcement learning problem.
@@ 3.2 Markov Decision Processes
In this chapter, we describe the general and formal model for reinforcement learning: Markov decision processes.
To ease the exposition, we start with a review of Markov Chains, also known as Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.

### v8

@@ 2.2.3 Quantitative Definitions of Stabilizability
The notion below is stronger in one sense and weaker in another: it allows adversarial bounded disturbances, but it asks only that the state remain uniformly bounded rather than converge to zero. For linear systems, related ideas include robust invariance and input-to-state stability.
@@
A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if the following holds. For every initial state satisfying $\|\mathbf{x}_{0}^{\pi}\|\leq 1$ and every disturbance sequence $\mathbf{w}_{0},\ldots,\mathbf{w}_{T-1}$ such that $\|\mathbf{w}_{t}\|\leq 1$ for all $t\in\{0,\ldots,T-1\}$, let $\mathbf{x}_{t}^{\pi}$ be the trajectory generated by policy $\pi$:
$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}^{\pi}),\mathbf{w}_{t}).$
Then
$\|\mathbf{x}_{t}^{\pi}\|\leq\gamma\qquad\text{for all }t=0,1,\ldots,T.$
@@
2.3 Complexity of Dynamical Systems and Hardness of Stabilizability
Before formally proving the computational hardness of stabilizability, we begin with two examples that illustrate how complicated invariant sets and attractors can become even for simple dynamical rules.
Consider the complex quadratic map
@@ 2.3 Complexity of Dynamical Systems and Hardness of Stabilizability
Its filled Julia set is
$K_{c}=\left\{z\in\mathbb{C}\ \colon\ \sup_{t\geq 0}|f_{c}^{t}(z)|<\infty\right\},$
where $f_{c}^{t}(z)$ denotes $t$ successive applications of the function $f_{c}$. The Julia set is the boundary $J_{c}=\partial K_{c}$. Depending on the parameter $c$, these sets can have extremely intricate geometry, and there are strong negative results on the computability of Julia sets for certain values of $c$; see, for example, [BY08]. A related parameter-space object is the Mandelbrot set, shown in Figure 2.6.
Another compelling visual illustration of complexity in dynamical systems is the Lorenz system, governed by the continuous-time equations
$\dot{x}$ $=\sigma(y-x),$
$\dot{y}$ $=x(\rho-z)-y,$
$\dot{z}$ $=xy-\beta z.$
For classical parameter values, trajectories are attracted to a complicated invariant set known as the Lorenz attractor, depicted in Figure 2.6. For background on this attractor, see [STR14, TUC99, LOR63]. These examples show that even describing the long-run behavior of a seemingly simple dynamical system can be a formidable task.
Figure 2.5: Mandelbrot set.
@@ 2.3.1 Computational Hardness of Stabilizability
Consider the following decision problem: given a dynamical system, can we determine whether it is asymptotically stabilizable, that is, whether there exists a control policy that drives the system to the zero state from every initial condition? We now show that this decision problem is computationally hard.
@@
For the family of discrete-time polynomial dynamical systems with integer coefficients, deciding asymptotic stabilizability is NP-hard.
@@
We reduce from 3SAT. Let
$F(x_{1:n})=c_{1}\land c_{2}\land\cdots\land c_{m}$
be a Boolean formula in conjunctive normal form with three literals per clause. Introduce real variables $u_{1},\ldots,u_{n}$. For each literal $\ell$, define
[…]
For a clause $c=\ell_{1}\lor\ell_{2}\lor\ell_{3}$, let
$g_{c}(\mathbf{u})=r(\ell_{1})\,r(\ell_{2})\,r(\ell_{3}).$
On Boolean assignments $\mathbf{u}\in\{0,1\}^{n}$, the quantity $g_{c}(\mathbf{u})$ equals $1$ exactly when the clause $c$ is false and 0 otherwise.
Now define the polynomial
$p_{F}(\mathbf{u})=\sum_{i=1}^{n}\bigl(u_{i}(1-u_{i})\bigr)^{2}+\sum_{j=1}^{m}g_{c_{j}}(\mathbf{u})^{2}.$
Then $p_{F}(\mathbf{u})=0$ if and only if $\mathbf{u}\in\{0,1\}^{n}$ encodes a satisfying assignment of $F$. Indeed, the first sum vanishes if and only if each $u_{i}\in\{0,1\}$, and the second sum vanishes if and only if no clause is false. Hence deciding whether a degree-six polynomial with integer coefficients has a real zero is NP-hard.
Consider now the scalar dynamical system
$x_{t+1}=p_{F}(\mathbf{u}_{t}).$
If $F$ is satisfiable, choose a satisfying assignment $\mathbf{u}^{\star}$ and apply the constant control $\mathbf{u}_{t}\equiv\mathbf{u}^{\star}$. Then $p_{F}(\mathbf{u}^{\star})=0$, so $x_{t}=0$ for all $t\geq 1$, regardless of the initial state. Therefore the system is asymptotically stabilizable.
Conversely, suppose $F$ is unsatisfiable. Then $p_{F}(\mathbf{u})>0$ for all $\mathbf{u}\in{\mathbb{R}}^{n}$. Since $p_{F}$ is continuous and
$\sum_{i=1}^{n}\bigl(u_{i}(1-u_{i})\bigr)^{2}\to\infty\qquad\text{as }\|\mathbf{u}\|\to\infty,$
the function $p_{F}$ attains a strictly positive global minimum $m>0$. Hence, for every control sequence,
$x_{t+1}=p_{F}(\mathbf{u}_{t})\geq m\qquad\text{for all }t\geq 0.$
So the state cannot converge to 0. Therefore the system is asymptotically stabilizable if and only if $F$ is satisfiable, proving NP-hardness. ∎
The proof already shows hardness for an especially simple family of polynomial control systems. Exercise 2.5 asks the reader to strengthen this result and show that deciding asymptotic stabilizability for polynomial systems of degree four is also NP-hard.
@@ 2.4 Bibliographic Remarks
The Lorenz equations were introduced to model the phenomenon of atmospheric convection [LOR63]. For detailed information on the mathematical properties of the Lorenz attractor, see [STR14, TUC99]. For discussion of computability questions for Julia sets and related sets, see [BY08].
Nonlinear dynamical systems admit many stability notions beyond those introduced here, including Lyapunov stability, local and global asymptotic stability, robust invariance, and input-to-state stability. Lyapunov’s direct method shows that suitable potential functions can certify stability without requiring explicit trajectory formulas; see [LYA92, KHA15] for further discussion.
An excellent survey on the computational complexity of control is [BT00]. For a more recent and stronger result on stabilizability, see [AHM12]. The NP-hardness theorem in this chapter appeared in [AHM16]. The computational complexity of mathematical programming—both constrained and unconstrained—is commented upon in [HAZ19].
@@ 2.5 Exercises
2.1. Strengthen Theorem 2.5 to show that deciding whether a dynamical system given by a polynomial of degree four is asymptotically stabilizable is NP-hard. Hint: Reduce from the NP-hard problem of MAX-2-SAT instead of 3-SAT.
@@
$\mathbf{x}_{t+1}=\begin{cases}\dfrac{A\mathbf{x}_{t}}{\|\mathbf{x}_{t}\|},&\mathbf{x}_{t}\neq\mathbf{0},\\[3.87498pt] \mathbf{0},&\mathbf{x}_{t}=\mathbf{0}.\end{cases}$
What are the equilibrium points $\mathbf{x}\in\mathcal{R}^{d}$ of the system? Which of them are stable, attractive, or asymptotically stable? Are any of them globally attractive? Your answer may depend on $A$.
2.3. Let $\alpha,\beta,\gamma>0$. Does the SIR model, as described in this chapter, define an asymptotically stabilizable dynamical system for a natural fixed choice of target state, such as a disease-free equilibrium? What about controllability? We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ and $\alpha u_{t}\leq S_{t}$ hold at all times.
@@ 2.4. Consider the double integrator system:
Assume throughout this exercise that $\mathbf{w}_{t}\equiv\mathbf{0}$.
@@
The previous chapters described and motivated the problem of control in dynamical systems. We have seen that, in full generality, even simple questions about dynamical systems such as stabilizability and controllability can be computationally hard. In this chapter, we study a more structured model for sequential decision making for which computationally efficient methods are available.
This chapter focuses on finite-state, finite-action, discounted, fully observed Markov decision processes. In this setting, the state space and transition dynamics are explicitly represented, and this extra structure makes efficient planning algorithms possible. In addition, throughout this chapter we assume that the transition probabilities and rewards are known. When these quantities are unknown and must be inferred from interaction with the environment, one arrives at reinforcement learning in the narrower sense.
After introducing the model of Markov decision processes, we discuss dynamic programming and the accompanying characterization of optimal policies via the Bellman equation. We then present two algorithmic consequences: a linear-programming formulation and the value-iteration algorithm.
@@ 3.1 Reinforcement Learning
Reinforcement learning is a subfield of machine learning that studies learning through interaction with a reactive environment. This is in contrast to learning from labeled examples, as in supervised learning, or inference in a probabilistic model, as in Bayesian statistics. In reinforcement learning, we have an agent and an environment. The agent observes its current state in the environment and takes an action; the environment then transitions to a new state and emits a reward.
Figure 3.1: The agent–environment interaction loop in reinforcement learning.
Typically, the goal is to learn behavior that maximizes the total reward accumulated over time. This viewpoint has led to some notable successes: machines have learned how to play Atari games and how to defeat the strongest human players at Go.
A crucial modeling assumption is the following.
The Markovian assumption: Conditional on the current state and current action, the distribution of the next state and immediate reward is independent of the past.
The reader can imagine situations where such an assumption is reasonable. For example, in the inverted pendulum setting from Section 2.1.3, the assumption is natural: once the current angle and angular velocity are known, the future evolution under a chosen torque does not depend on how that state was reached.
Other problems call this assumption into question. In the medical ventilation problem, for instance, future treatment may depend on the medical history of the patient and not only on the present physiological variables. One possible remedy is to expand the state description so that it includes enough information to make the process Markovian. In the extreme, the state can contain the entire history, in which case the Markovian assumption becomes true by construction. Of course, such a representation may be computationally inefficient.
In this chapter we restrict attention to the fully observed setting, in which the state itself is available to the agent. When only partial observations are available, the appropriate formal model is a partially observed Markov decision process (POMDP), which is beyond our present scope.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom-left square, and its goal is to reach the terminal state with reward +1+1. The square marked 1-1 is also terminal, but undesirable. The block ‘X’ represents an obstacle through which the robot cannot travel.
@@
The robot can choose to move left, right, up, or down. However, the controls are noisy. With probability $0.8$, the robot moves in the intended direction; with the remaining probability $0.2$, it moves sideways, with probability $0.1$ to each side.
Notice that the Markovian assumption clearly holds in this example. Once the robot’s current square and chosen action are specified, the distribution of the next square and the immediate reward no longer depend on the earlier history.
A solution to this problem is a mapping from states to actions. Such a mapping is called a policy. The best policy depends not only on the transition structure but also on how rewards are assigned. To connect the example to the discounted infinite-horizon formalism used later in the chapter, it is convenient to model terminal states as absorbing states: entering the upper terminal state yields reward +1+1, entering the lower terminal state yields reward 1-1, and thereafter the process remains in that terminal state with reward 0.
For example, if the reward for each non-terminal step is mildly negative, say 0.04-0.04, then Bubbo prefers to steer toward the +1+1 terminal while avoiding the risk of falling into the 1-1 terminal, as in Figure 3.3.
@@
Figure 3.3: An optimal policy when the reward for each non-terminal step is 0.04-0.04.
Conversely, if the step cost is very negative, then Bubbo prefers to terminate the episode as quickly as possible, even at the risk of ending in the 1-1 terminal. This leads to a different optimal policy, as in Figure 3.4.
@@
Figure 3.4: If the reward of each non-terminal step is changed to 2-2, the optimal policy changes: the robot is now more concerned with ending the game quickly than with reaching the +1+1 terminal.
We next describe a formal model that captures this and many other reinforcement-learning problems.
@@ 3.2 Markov Decision Processes
We now describe the mathematical model used in this chapter: finite discounted Markov decision processes. To ease the exposition, we begin with Markov chains, proceed to Markov reward processes, and then give the full Markov decision process definition.